# Auditing and Generating Synthetic Data with Controllable Trust Trade-offs

### v3

@@ 2 Figure and Table Captions
Caption of Table 3 Performance of ResNet50 on ImageNet v1 and v2 Datasets under varied noisy conditions (additive random Gaussian noise to tested images) and different training regimes (synthetic data augmentation). The resulting models are evaluated in Acc@1 (first number) and Acc@5 (second number). We see that while classifiers trained purely on synthetic data lag behind those trained on real, augmenting real data with synthetic images makes models more resilient to noise.
@@ Acknowledgments
where the downstream task and auditing configurations cfg are defined as follows:
@@
In Extended Data Table 4, we provide the set of metrics corresponding to each trust dimension and their associated polarities. Note that all Fidelity and Privacy metrics are evaluated between the synthetic data $D_{s}$ and the real training set $D_{r,\text{train}}$. Utility, Fairness, and Robustness metrics refer to evaluations of classifiers trained on the synthetic data $D_{s}$, validated on the real development set $D_{r,\text{val}}$, with final measurements reported on the real test set $D_{r,\text{test}}$. A glossary of all metrics and their definitions is provided in Supplementary Information C.
@@
    Empirical Cumulative Distribution Function (ECDF) evaluation: For each aligned metric $\tilde{m}_{i,T},$ we compute the Empirical CDF, given the observations across multiple synthetic datasets $D^{j}_{s},\leavevmode\nobreak\ j=1,\ldots,N$:
@@
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ u_{T,i}=\text{ECDF}_{\tilde{m}_{i,T}}(\tilde{m}_{T,i}(D_{r},D_{s},\text{{cfg}})),\leavevmode\nobreak\ i=1,\dots,M_{T}.$ (5)
@@
    $\centering\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \pi_{T}(D_{r},D_{s},\text{{cfg}})=\exp\left(\frac{1}{M_{T}}\sum_{i=1}^{M_{T}}\log(u_{i,T})\right).\@add@centering$ (6)
@@
    $\centering\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \pi_{T}(D_{r},D_{s},\text{{cfg}})=\exp\left(\sum_{i=1}^{M_{T}}\beta_{i,T}\log(u_{i,T})\right).\@add@centering$ (7)
The details of the implementation of the copula aggregation method are given in Supplementary Information D. Although effective and widely accepted as an aggregation method, the main limitation of the geometric mean copula is the inherent assumption of independence between metrics. This shortcoming can be alleviated if a dependency structure is known and can be incorporated via hierarchical or vine copulas [97].
@@
We are now ready for the last steps in our auditing framework Reweight & Index by Trust. Trust dimensions can often be in conflict with each other, hence, for each application and use case and depending on the policy and regulation landscape, a trade off between these dimensions can be defined by introducing a weighing scheme $\omega$ for each trust dimension. This weighting scheme reflects the order of priorities among trust dimensions to ensure compliance with safeguards and regulations. The weights
@@
define a probability distribution on trust dimensions i.e $\omega_{T}\geq 0,\sum_{T\in\mathcal{T}}\omega_{T}=1.$ Armed with these trade off weights and the trust dimension indices, we define the trustworthiness index of a synthetic dataset $D_{s}$ as the weighted geometric mean of all trust dimension indices:
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \tau_{\text{Trust}}(D_{r},D_{s},\text{{cfg}},\omega)=\exp\left(\sum_{T\in\mathcal{T}}\omega_{T}\log(\pi_{T}(D_{r},D_{s},\text{{cfg}}))\right),$ (8)
where $\pi_{T}$ are trust dimension indices defined in Equations (6) or (7). From now on , we use the notation $\omega=(\omega_{f},\omega_{P},\omega_{U},\omega_{F},\omega_{R})$ to refer to trust dimensions trade-offs corresponding to Fidelity, Privacy, Utility, Fairness, and Robustness, respectively. Table 2 provides examples of weighting schemes corresponding to different priorities in auditing synthetic data.
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ R^{\alpha}_{\text{trust}}(\mathcal{D}_{r},\mathcal{D}^{j}_{s},\text{{cfg}},\omega)=\log(\overline{\tau_{\text{Trust}}}(\mathcal{D}_{r},\mathcal{D}^{j}_{s},\text{{cfg}},\omega))-\alpha\log(\Delta_{\tau}(\mathcal{D}_{r},\mathcal{D}^{j}_{s},\text{{cfg}},\omega)),\leavevmode\nobreak\ \leavevmode\nobreak\ j=1\dots N.$ (12)
@@
Our auditing framework sets templates for communicating the audit results for all synthetic data in the form of an auditing report. The auditing report starts by giving transparency about the real data used to train the generative models, and the training and inference configurations used for training and sampling synthetic data from each generative model. Then, for a given trade off weighting of trust dimensions $\omega$, a ranked list of models is produced using the trustworthiness index $\tau_{trust}$ defined in Equation (8) or using ranking under uncertainty of trust indices given in Equations (11) or (12). Synthetic data cards in the ranked list contain the overall trustworthiness index $\tau_{\text{trust}}$, as well as the indices for each trust dimension along their variances. In addition, each card is paired with an interpretable message and warnings about potential harms and trustworthiness violations detected in the synthetic data for each trust dimension. Finally, a breakdown of all metrics evaluated for each dimension is presented, comparing all synthetic datasets to provide further insights on the aggregated trustworthiness index. An example of the audit report is given in Supplementary Information G.
@@
$\log\left(\frac{\mathbb{P}(A(D)\in\mathcal{O})}{\mathbb{P}(A(D^{\prime})\in\mathcal{O})}\right)\leq\varepsilon,\forall\leavevmode\nobreak\ \text{neighboring }D,D^{\prime}.$
@@
$\pi_{\text{Fidelity}}(D_{r},D^{t}_{s},\text{{cfg}}),\leavevmode\nobreak\ \pi_{\text{Privacy}}(D_{r},D^{t}_{s},\text{{cfg}}),\leavevmode\nobreak\ \pi^{\text{val}}_{\text{Utility }}(D_{r},D^{t}_{s},\text{{cfg}}),\leavevmode\nobreak\ \pi^{\text{val}}_{\text{Fairness }}(D_{r},D^{t}_{s},\text{{cfg}}),\leavevmode\nobreak\ \pi^{\text{val}}_{\text{Robustness }}(D_{r},D^{t}_{s},\text{{cfg}}),$ (13)
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \tau_{\text{trust}}^{val}(D_{r},D^{t}_{s},\text{{cfg}},\omega)$ $=\omega_{f}\pi_{\text{Fidelity}}(D_{r},D^{t}_{s},\text{{cfg}})+\omega_{P}\pi_{\text{Privacy}}(D_{r},D^{t}_{s},\text{{cfg}})+\omega_{U}\pi^{\text{val}}_{\text{utility }}(D_{r},D^{t}_{s},\text{{cfg}})$
$+\omega_{F}\pi^{\text{val}}_{\text{Fairness }}(D_{r},D^{t}_{s},\text{{cfg}})+\omega_{R}\pi^{\text{val}}_{\text{Robustness }}(D_{r},D^{t}_{s},\text{{cfg}})$ (14)
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ t^{*}=\arg\max_{t\in\{1,\dots I_{\max}\}}\tau_{\text{trust}}^{val}(D_{r},D^{t}_{s},\text{{cfg}},\omega),$ (15)

### v4

@@ Methods
TABLE IX: Performance of ResNet50 on ImageNet v1 and v2 Datasets under varied noisy conditions (additive random Gaussian noise to tested images) and different training regimes (synthetic data augmentation). The resulting models are evaluated in Acc@1 (first number) and Acc@5 (second number). We see that while classifiers trained purely on synthetic data lag behind those trained on real, augmenting real data with synthetic images makes models more resilient to noise.
Synthetic data generation
We generated several privacy-preserving synthetic datasets using our non-private pretrained TabGPT model. For model selection in this experiment we relied on a fidelity validation of the TabGPT model. To generate private synthetic data, we used a private sampling technique [92] , which involves adding Laplacian noise with controlled variance (dependent on the user-provided $\varepsilon$ value) to the probability distribution over the generated tokens from the non-private GPT model. This is a form of output-perturbation methods that guarantees differential privacy.We generated seven datasets with varying privacy levels, from highly private ($\varepsilon=0.1$) to non-private ($\varepsilon=200$), as shown in the rows labeled "Synthetic" in Table VII. Additionally, we considered real card transaction data and synthetically generated data without private sampling.
Training the downstream Fraud Detection Model
Results and Discussion
We audit the synthetic dataset sampled from different epochs (namely after 3, 5, 7 and 9) during the fine-tuning process. The downstream task we consider in this audit is the in hospital mortality prediction task, where the protected community is the ethnicity "ASIAN" [28]. As an embedding E, we use the original pre-trained BioGPT model [71] to extract embeddings for the synthetic notes as it has the capability of representing the biomedical domain. In Table VIII we see that our trustworthiness index driven model selection allows a controllable trade-off between utility and fairness : epoch 7 has a better utility fairness trade-off than the model at the last epoch that has higher utility. Therefore it is favorable to select the model at epoch 7 at the price of a reduced utility but with an enhanced fairness, which is of paramount importance for this use-case.
We consider here the use of synthetic image data for training classifiers on one of the key computer vision datasets, Imagenet [35], with a focus on the utility and generalization propreties of these classifiers on the Imagenet test set and under distribution shifts. Imagenet consists of $1.2$ Million images of size 256x256x3 and 1000 categories. Authors in [93] constructed Imagenetv2 test set, that consists of a distribution shift from the original imagenet distribution and reported a performance drop between $11\%-14\%$. Recent works [12] and [22] showed promising results of synthetic data from diffusion models (a family of generative models that uses diffusion techniques [100]) in improving Imagenet classification. Following these promising works, we synthesize $1.2$ M labeled images from the Imagenet 256x256 pretrained guided diffusion models from OpenAI [36].Table IX presents ResNet50 [50] performance on ImageNet-v1 and v2 under Gaussian noise, considering four training data scenarios: (1) real Imagenet; (2) synthetic data; (3) a hybrid with real and synthetic images (real to synthetic ratio $1/0.5$); and (4) an equal mix totaling 2.4 million images. Results show synthetic-only models lag, but integrating with real images enhances robustness, especially in noisier settings (sigma 0.3 and 0.5). Combining synthetic data with real images improves a model’s resilience to noise and image corruption.
Conclusion
Figure 6: Iterative Use of Synthetic Data to Train Generative Models on the Bank Marketing Dataset. The plot shows mean and variance of indices across folds for all generations. As we iteratively use synthetic data to train generative models, we see that while the fidelity index of the synthetic datasets collapses, the privacy index improves. The utility of the synthetic data on downstream tasks fluctuates but is essentially constant. Note that the utility of these models does not deteriorate, since the index of these synthetic data at first iteration is not high.
Ablation Studies
Debiasing downstream tasks trained on Synthetic Data
We notice on tabular datasets that while private synthetic data with high performing robustness index are still competitive in terms of their utility, the ones with high fairness index are not competitive in terms of their utility. In order to further improve the fairness, privacy and utility tradeoffs, we explore including debiased downstream classifiers in Supplementary Information Tables VI, X and -Y. Debiasing leads to an improvement fairness while maintaining utility of private synthetic data.
Independent Copula Versus Dependent Copula
In our aggregation we used an independent copula as a mean to aggregate the metrics. We explore here using a Gaussian copula to capture the dependencies between metrics and see how it impacts the overall ranking of the synthetic datasets on all tabular datasets we consider. We explore Gaussian copulas, since estimating dependent copulas from data is cursed in high dimension. To estimate the Gaussian copula, we compute the mean and covariance matrix for the concatenation of all metrics under a trust dimension, and compute the multivariate Gaussian CDF. This comes with two caveats: 1) the dependency may not be Gaussian 2) given small number of samples the covariance matrix is not well estimated and is low rank and inverting it in the computation of the multivariate CDF leads to numerical instabilities that we circumvent with a regularization of the covariance matrix $\Sigma$ by $\Sigma+\lambda I$, for $\lambda=10^{-3}$. We first report the average time in seconds to compute the independent copula on tabular use cases on an Intel (R) Xeon (R) CPU E5-2667 v2 @ 3.30GHz that is of $11s$ versus $1700s$ for the Gaussian copula (See Supplementary Information -R Table 3). Second we report the overlap between the ranked list of synthetic data returned by the independent copula $L_{IC}$ and the one returned by the Gaussian copula $L_{GC}$:
[…]
From Supplementary Information -R Figure 4 we see that there is 60-70% overlap between the top-10 ranking synthetic datasets returned by independent and Gaussian copula. In light of these results, given the simplicity and the speed of the independent copula and the numerous issues of the Gaussian copula in terms of its slow computation and bad conditioning and unfavorable statistical guarantees, the independent copula is preferable and offers a good computational tradeoff in estimating the copula.
Synthetic Data Model Gaussian Copula CTGAN PateGAN
$1-\textrm{AUC}$ of Membership Attack [108] 0.4966 0.4882 0.4879
Privacy Index 0.2916 0.2866 0.2588
Table X: Membership Inference Attacks of [108] versus our Privacy Index on the Bank Marketing dataset. In the second row we give $1-\textrm{AUC}$ of the membership attack classifier (where AUC is the Area Under the Curve). Higher values of $1-\textrm{AUC}$ means that the membership attack at identifying members of the training data is less successful, meaning higher privacy. We see that this membership attack lead to the same ranking of synthetic data as our privacy index given in the third row.
Membership attacks
The membership attacks we used in our privacy auditing were based on nearest neighbor attacks that assess memorization issues in synthetic data. Identity disclosure attacks assess privacy in synthetic data beyond memorization [104],[39]. We explore here the impact of other targeted membership attacks based on likelihood ratios between synthetic data and a reference to estimate the membership [108]. This attack is the state of the art membership inference method. We train this attack on the Bank Marketing dataset, evaluating synthetic data of three baselines: the Gaussian copula, CTGAN and PateGAN. We give in in ] Table X, $1-\textrm{AUC}$ of this attack, where the Area Under the Curve (AUC) is evaluated on the real data with members and non members of the training set. Higher values of $1-\textrm{AUC}$ mean that the membership attack is less successful at identifying members of the real data, meaning higher privacy. We see in Table X that this membership attack lead to the same ranking of synthetic datasets as our privacy index (computed without the inclusion of the membership attack). Those membership inference attacks can be seamlessly added to our privacy index computation.
Iterative Collapse Risk of Synthetic Data
As synthetic data becomes more and more prevalent, the risks of iteratively training generative models on synthetic data becomes more and more pressing. Authors in [99] referred to this issue as the iterative curse or iterative collapse in synthetic data. They also showed that the fidelity of the synthetic data deteriorates as we iteratively train generative models on synthetic data. We show in Figure 6 on the Bank Makerting dataset, the audits of synthetic data obtained from training models successively on synthetic data generated from the previous iteration. As we iteratively use synthetic data to train generative models, we see in Figure 6 that while the fidelity index of the synthetic datasets collapses, the privacy index improves as expected. The utility of the synthetic data on downstream tasks fluctuates but is essentially constant. Note that the utility of these models does not deteriorate, since the utility of these synthetic data at first iteration is not high.
Data availability
Data used in the paper is available online and open source. See Table III for references.
Code availability
@@ Acknowledgments
Brian Belgodere is a senior research software engineer in the Emerging Tech. Engineering Department at IBM Research. He currently works on hybrid cloud platforms and tooling for Machine Learning and AI. He holds B.S. degrees in economics and business administration from Carnegie Mellon University and a Juris Doctor from the University of Pittsburgh.
Pierre Dognin received his M.S. degree and Ph.D degree in Electrical Engineering from the University of Pittsburgh in 1999 and 2003. He then joined IBM Research in the Human Language Technologies Department where he worked as a Research Staff Member on Automatic Speech Recognition, focusing on acoustic modeling, robust speech recognition, and core algorithm technology. Since 2015, he has been working on multimodal Machine Learning and is now in the Trusted AI Department where his research interests include Deep Learning, Generative Modeling for Computer Vision and Natural Language.
Igor Melnyk is currently a Research Staff Member in the Trusted AI Department at IBM Research. Before that he completed his PhD degree in Computer Science and Engineering in 2016 from the University of Minnesota. In 2010, he obtained his MS degree in Computer Science from the University of Colorado, Boulder. His research interests are in the areas of Machine Learning and AI with the focus on the problems of Natural Language Processing and Computer Vision.
Youssef Mroueh is a research staff member in the Trusted AI department at IBM Research and a principal investigator in the MIT-IBM Watson AI lab. He received his PhD in computer science in February 2015 from MIT, CSAIL, where he was advised by Professor Tomaso Poggio. In 2011, he obtained his engineering diploma from Ecole Polytechnique Paris France, and a master of science in Applied Maths from Ecole des Mines de Paris. He is interested in Multimodal Deep Learning, Generative modeling, Computer Vision and Learning Theory.
Aleksandra (Saška) Mojsilović is a Serbian-American scientist. Her research interests are artificial intelligence, data science, and signal processing. She is known for innovative applications of machine learning to diverse societal and business problems. Her current research focuses on issues of fairness, accountability, transparency, and ethics in AI. She is an IBM Fellow and IEEE Fellow.
Jiri Navratil received M.Sc. (’95) in EE & PhD (’99) in CS from Technische Universitaet Ilmenau, Germany. He joined IBM Research in 1999, where he is currently a Principal Research Staff Member. His research interests include Deep Learning, Natural Language Processing and Industrial AI Applications.
Apoorva Nitsure is a Research Software Engineer at IBM Research. Her interests include Deep Learning, Natural Language Processing and Interpretability in Machine Learning to help users make informed choices while using AI. Prior to joining IBM Research, she completed her Masters degree from Heinz College, Carnegie Mellon University and B.Tech in IT from CoEP India.
Inkit Padhi received Masters’s degree in Computer Science from Viterbi School of Engineering, USC/ISI in May 2016. Since December of 2018, he is working as a research engineer in the Trusted AI department at IBM Research. He is interested in Machine Learning and Natural Language Processing. His pronouns are he, him, and his.
Mattia Rigotti is a Research Staff Member in the AI Automation Department at IBM Research, which he joined in 2014. He holds an M.Sc. degree in Theoretical Physics from ETH Zurich and a Ph.D. in Computational Neuroscience from Columbia University. His research interests include Deep Learning, Neuromorphic Engineering and Computational Neuroscience.
Jarret Ross received an M.S. in Computer Science from Wichita State University. He has been with IBM research since 2016 and he is currently a research engineer in the Trusted AI department. His interests are in deep learning, distributed computing and generative modeling.
Yair Schiff is a PhD student in the Computer Science department at Cornell University. Yair was a software developer working on IBM Watson Studio. Prior to joining IBM, he completed an M.S. degree in Computer Science at the Courant Institute of Mathematical Sciences at New York University. Yair collaborated with the Trusted AI department at IBM Research. .
Richard A. Young received his B.S. Hons degree in computer science from The University of the Witwatersrand (South Africa) in 2010. He worked as a software developer in the banking and health insurance industries from 2011 to 2018. He joined IBM Research in South Africa as a Research Engineer in 2018 where he specialises in building web and cloud based applications.
Radhika Vedpathak is a designer with an experience in developing design for web applications.
Richard A. Young received his B.S. Hons degree in computer science from The University of the Witwatersrand (South Africa) in 2010. He worked as a software developer in the banking and health insurance industries from 2011 to 2018. He joined IBM Research in South Africa as a Research Engineer in 2018 where he specialises in building web and cloud based applications.
@@
In Table I, we provide the set of metrics corresponding to each trust dimension and their associated polarities. Note that all Fidelity and Privacy metrics are evaluated between the synthetic data $D_{s}$ and the real training set $D_{r,\text{train}}$. Utility, Fairness, and Robustness metrics refer to evaluations of classifiers trained on the synthetic data $D_{s}$, validated on the real development set $D_{r,\text{val}}$, with final measurements reported on the real test set $D_{r,\text{test}}$. A glossary of all metrics and their definitions is provided in Supplementary Information -K.
@@
    Empirical Cumulative Distribution Function (ECDF) evaluation: For each aligned metric $\tilde{m}_{i,T},$ we compute the Empirical CDF, given the observations across multiple synthetic datasets $D^{j}_{s},~{}j=1,\ldots,N$:
@@
    $u_{T,i}=\text{ECDF}_{\tilde{m}_{i,T}}(\tilde{m}_{T,i}(D_{r},D_{s},\text{{cfg}})),~{}i=1,\dots,M_{T}.$ (4)
@@
    $\centering\pi_{T}(D_{r},D_{s},\text{{cfg}})=\exp\left(\frac{1}{M_{T}}\sum_{i=1}^{M_{T}}\log(u_{i,T})\right).\@add@centering$ (5)
@@
    $\centering\pi_{T}(D_{r},D_{s},\text{{cfg}})=\exp\left(\sum_{i=1}^{M_{T}}\beta_{i,T}\log(u_{i,T})\right).\@add@centering$ (6)
The details of the implementation of the copula aggregation method are given in Supplementary Information -Q. Although effective and widely accepted as an aggregation method, the main limitation of the geometric mean copula is the inherent assumption of independence between metrics. This shortcoming can be alleviated if a dependency structure is known and can be incorporated via hierarchical or vine copulas [33].
@@
$\tau_{\text{Trust}}(D_{r},D_{s},\text{{cfg}},\omega)=\exp\left(\sum_{T\in\mathcal{T}}\omega_{T}\log(\pi_{T}(D_{r},D_{s},\text{{cfg}}))\right),$ (7)
where $\pi_{T}$ are trust dimension indices defined in Equations (5) or (6). From now on , we use the notation $\omega=(\omega_{f},\omega_{P},\omega_{U},\omega_{F},\omega_{R})$ to refer to trust dimensions trade-offs corresponding to Fidelity, Privacy, Utility, Fairness, and Robustness, respectively. Table II provides examples of weighting schemes corresponding to different priorities in auditing synthetic data.
@@
${}R^{\alpha}_{\text{trust}}(\mathcal{D}_{r},\mathcal{D}^{j}_{s},\text{{cfg}},\omega)=\log(\overline{\tau_{\text{Trust}}}(\mathcal{D}_{r},\mathcal{D}^{j}_{s},\text{{cfg}},\omega))-\alpha\log(\Delta_{\tau}(\mathcal{D}_{r},\mathcal{D}^{j}_{s},\text{{cfg}},\omega)),~{}~{}j=1\dots N.$ (11)
@@
Our auditing framework sets templates for communicating the audit results for all synthetic data in the form of an auditing report. The auditing report starts by giving transparency about the real data used to train the generative models, and the training and inference configurations used for training and sampling synthetic data from each generative model. Then, for a given trade off weighting of trust dimensions $\omega$, a ranked list of models is produced using the trustworthiness index $\tau_{trust}$ defined in Equation (7) or using ranking under uncertainty of trust indices given in Equations (10) or (11). Synthetic data cards in the ranked list contain the overall trustworthiness index $\tau_{\text{trust}}$, as well as the indices for each trust dimension along their variances. In addition, each card is paired with an interpretable message and warnings about potential harms and trustworthiness violations detected in the synthetic data for each trust dimension. Finally, a breakdown of all metrics evaluated for each dimension is presented, comparing all synthetic datasets to provide further insights on the aggregated trustworthiness index. An example of the audit report is given in Supplementary Information -T.
@@
$\log\left(\frac{\mathbb{P}(A(D)\in\mathcal{O})}{\mathbb{P}(A(D^{\prime})\in\mathcal{O})}\right)\leq\varepsilon,\forall~{}\text{neighboring }D,D^{\prime}.$
@@
$\pi_{\text{Fidelity}}(D_{r},D^{t}_{s},\text{{cfg}}),~{}\pi_{\text{Privacy}}(D_{r},D^{t}_{s},\text{{cfg}}),~{}\pi^{\text{val}}_{\text{Utility }}(D_{r},D^{t}_{s},\text{{cfg}}),~{}\pi^{\text{val}}_{\text{Fairness }}(D_{r},D^{t}_{s},\text{{cfg}}),~{}\pi^{\text{val}}_{\text{Robustness }}(D_{r},D^{t}_{s},\text{{cfg}}),$ (12)
@@
$\tau_{\text{trust}}^{val}(D_{r},D^{t}_{s},\text{{cfg}},\omega)$ $=\omega_{f}~{}\pi_{\text{Fidelity}}(D_{r},D^{t}_{s},\text{{cfg}})$
$+\omega_{P}~{}\pi_{\text{Privacy}}(D_{r},D^{t}_{s},\text{{cfg}})+\omega_{U}~{}\pi^{\text{val}}_{\text{utility }}(D_{r},D^{t}_{s},\text{{cfg}})$
$+\omega_{F}~{}\pi^{\text{val}}_{\text{Fairness }}(D_{r},D^{t}_{s},\text{{cfg}})+\omega_{R}~{}\pi^{\text{val}}_{\text{Robustness }}(D_{r},D^{t}_{s},\text{{cfg}})$ (13)
@@
$t^{*}=\arg\max_{t\in\{1,\ldots,I_{\max}\}}\tau_{\text{trust}}^{val}(D_{r},D^{t}_{s},\text{{cfg}},\omega),$ (14)